OpenStudy (anonymous):

let theta be an angle in quadrant IV such that costheta=4/5 find the exact values of csc theta and tan theta

OpenStudy (anonymous):

\[\cos \theta=\frac{ 4 }{ 5 }\] find \[\csc \theta \] and \[\tan\theta\]

OpenStudy (campbell_st):

well draw a diagram  hope it helps